# The Galaxy Activity, Torus, and Outflow Survey (GATOS). (VI): Polycyclic Aromatic Hydrocarbon Emission in the Central Regions of Three Seyferts

### v1

@@ 1 Introduction
Leveraging capabilities of the Medium Resolution Spectrograph (MRS; Wells et al. 2015; Labiano et al. 2021; Argyriou et al. 2023) on the Mid-Infrared Instrument (MIRI; Rieke et al. 2015; Wright et al. 2015, 2023) onboard JWST, this manuscript showcases the diversity of PAH characteristics within the central kpc scale regions of three nearby Seyfert galaxies. The goal of this manuscript is combining the analysis of PAH features (Section 3) and other emission line diagnostics (Section 4), to shed light on the underlying mechanisms responsible for the varying PAH characteristics of the three targets (Section 5).
@@ 2.1 Targets and Analysis
This manuscript is part of a series studying in total of six type 1.9/2 Seyferts with JWST MIRI/MRS integral field unit (IFU) spectral observations obtained by the JWST cycle 1 GO program (#1670; PI: Shimizu, T. Taro). The six targets are all included in the Galactic Activity, Torus, and Outflow Survey (GATOS; García-Burillo et al. 2021, 2024; Alonso-Herrero et al. 2021; García-Bernete et al. 2024a; Zhang et al. 2024)111https://gatos.myportfolio.com. The full GATOS sample is a nearly complete selection of AGN with luminosities at distances of Mpc from the 70 Month Swift-BAT All sky Hard X-ray Survey (Baumgartner et al. 2013). This sample is also largely unbiased to obscuration/absorption even up to column densities of , from which the six targets of GO program #1670 are further selected based on their outflow properties. A first analysis of the data set of this program, focussed on silicate features and water ice at small scales, is presented by García-Bernete et al. (2024a), followed by analyses focused on different aspects of one or more of the six targets (e.g., Davies et al. 2024; Esparza-Arredondo et al. 2024; García-Bernete et al. 2024c; Hermosa Muñoz et al. 2024; Zhang et al. 2024, and Delany et al. in prep., Haidar et al. in prep.).
@@
In this manuscript, we focus on the remaining three targets, which exhibit relatively weak PAH emission without significantly extended outflow regions and cover a considerable range of AGN strength. The goal of this work is to shed light on the specific mechanisms that could be responsible for the variation of PAH properties around different AGN. To this end, this manuscript combines the analysis of both observational measurements and theoretical models of not only PAH features but also other infrared emission lines, which is among the first attempts of such work in the era of JWST. Given the relatively weak PAH emission of the three targets studied here, after some testing work, we choose to proceed the analysis with the MRS spectra extracted with a series of apertures (see Figure 1, 500 pc in physical scale for each, well beyond the nuclear regions studied by García-Bernete et al. 2024c), rather than the spatially resolved MRS spectra. These apertures are simply selected to match the field of view of MRS IFU units.
@@ 3 PAH Characteristics of Central Apertures in Target Seyferts
Compared to the SF spaxels, the apertures of ESO137-G034 overall exhibit the much larger PAH 11.3 /7.7 band ratios and slightly larger PAH 6.2 /7.7 band ratios, while the aperture of MCG-05-23-016 exhibits the much smaller PAH 6.2 /7.7 band ratio and the moderate PAH 11.3 /7.7 band ratio. Moreover, the apertures of ESO137-G034 and MCG-05-2-016 on the PAH diagram are close to the loci of low metallicity star-forming systems (i.e., H II regions and BCDs) and low redshift quasars, respectively. We will specifically discuss the similarity and difference between these systems in Section 4 to shed more light on the physical mechanisms that could be responsible for their PAH characteristics. Meanwhile, the apertures of NGC 3081 exhibit the PAH band ratios of a wide distribution, overlapping with those of a large sample of Seyfert galaxies studied in the era of Spitzer. Overall, these apertures have a wide distribution of PAH band ratios, covering those of most Seyferts, as well as those of the outflow regions and the innermost regions of the three targets studied in Paper I.
@@ 4 Emission Line Diagnostics of Central Apertures in Target Seyferts
We find that although the three apertures in ESO137-G034 exhibit the ionized emission line ratios consistent with the results obtained by AGN models, their highest ionization potential lines (e.g., [Ne V] with the ionization potential, i.e., IP, of 97.1 eV) are remarkably weak compared to other high ionization potential lines (e.g., [O IV] and [S IV] with the IP of 54.9 eV and 34.8 eV, respectively). For reference, the apertures in NGC 3081, which are also dominated by AGN excitation, exhibit overall strong emission of all the three high ionization potential lines. Specifically, AGN models calculated with incident radiation SEDs of lower Eddington ratios and smaller ionization parameters, i.e., as the indicator of ionization field intensity, exhibit lower [Ne V]/[Ne II] ratios (Pereira-Santaella et al. 2024). Accordingly, the lower [Ne V]/[Ne II] ratios reflect the weak accretion nature and moderately intense radiation field of ESO137-G034. Meanwhile, the radiation filed therein is hard given the large values of the [Ne III]/[Ne II] ratio, which is a canonical indicator of the radiation filed hardness (e.g., Thornley et al. 2000; Groves et al. 2006). Additionally, we find that the central region of MCG-05-23-016 is likely also affected by fast radiative shocks in addition to AGN excitation. However, the emission lines from the central aperture of MCG-05-2-016 do not show the significant broadening as expected for highly shocked regions. This suggests that the fast radiative shocks in MCG-05-2-016 are more likely playing a role in the radiative rather than kinetic manner as detailed later.
@@
Related to this result, we find that [Fe II]/Pf and [Fe II]/Hu ratios, as another empirical diagnostics of collisional shock heating, of the three targets exhibit the same trend as their ratios. Namely, the apertures of ESO137-G034 and MCG-05-23-016 exhibit the highest and lowest [Fe II]/Pf as well as [Fe II]/Hu ratios, respectively, while the apertures of NGC 3081 exhibits the values between those of ESO137-G034 and MCG-05-23-016 (see Figure A3). In fact, the intrinsic [Fe II] emission of MCG-05-23-016 is not weak, and its value is roughly the same as that of the innermost aperture of NGC 3081 and greater than those of other apertures of NGC 3081 (see Table A2). The aperture of MCG-05-23-016 exhibits the relatively weak evidence of collisional shock heating, but have the strongest [Ne II] emission and the highest Pf/ and Hu/ ratios among all the 13 apertures of the three targets (see Figure A2). Meanwhile, the three apertures of ESO137-G034 have the highest temperatures according to their ratios, consistent with its strongest signature of collisional shock heating. These results, combining the findings from Figure 4, indicate that the central region of MCG-05-23-016 is more likely dominated by photo-ionization from the (fast radiative) shock precursor in addition to the AGN, i.e., the radiative mode feedback, rather than the collisional shock excitation, i.e., the kinetic mode feedback, while the central region of ESO137-034 is highly heated due to the collisional shock excitation (see further discussion in Section 5).
@@ 5 Discussion: Underlying Mechanisms Responsible for the PAH Characteristics
Accordingly, combining the PAH characteristics and other emission line diagnostics covered by JWST spectra, we are able to have a more comprehensive understanding of different feedback effects of AGN activity, which is pivotal for fully understanding the galaxy formation and evolution. Further study, similar to the analysis done here, of a larger sample and for different AGN systems (e.g., quasars, LINERs) could allow us to use PAH characteristics as diagnostics of different evolutionary processes in galaxies, which has a promising application in the era of JWST.
@@ 6 Summary and Conclusions
Leveraging high quality JWST MIRI/MRS IFU observations, this manuscript showcases the distinct PAH characteristics in a sample of three Seyferts, based on the PAH measurements from a series of (i.e., 500 pc) apertures (Section 2). We find that positions occupied by apertures in the three targets on the PAH diagram are different from those occupied by regions in star-forming galaxies. Specifically, we find larger PAH 11.3 /7.7 ratios or smaller PAH 6.2 /7.7 ratios for apertures in the three targets, indicating overall more neutral or larger PAHs therein compared to regions in star-forming galaxies (Section 3). In addition to the PAH diagram, we also present and discuss other emission line diagnostics, revealing the existence of extra heating or potential fast radiative shocks in these targets (Section 4). Combining the PAH diagram and the emission line diagnostics, we further discuss the underlying mechanisms responsible for the PAH characteristics of the three targets (Section 5).

### v2

# Polycyclic Aromatic Hydrocarbon Emission in the Central Regions of Three Seyferts and the Implication for Underlying Feedback Mechanisms

@@ 1 Introduction
Leveraging capabilities of the Medium Resolution Spectrograph (MRS; Wells et al. 2015; Labiano et al. 2021; Argyriou et al. 2023) on the Mid-Infrared Instrument (MIRI; Rieke et al. 2015; Wright et al. 2015, 2023) onboard JWST, this letter showcases the diversity of PAH characteristics within the central kpc scale regions of three nearby Seyfert galaxies. The goal of this letter is combining the analysis of PAH features (Section 3) and other emission line diagnostics (Section 4), to shed light on the underlying mechanisms responsible for the varying PAH characteristics of the three targets (Section 5).
@@ 2.1 Targets and Analysis
This letter is part of a series studying in total of six type 1.9/2 Seyferts with JWST MIRI/MRS integral field unit (IFU) spectral observations obtained by the JWST cycle 1 GO program (#1670; PI: Shimizu, T. Taro). The six targets are all included in the Galactic Activity, Torus, and Outflow Survey (GATOS; García-Burillo et al. 2021, 2024; Alonso-Herrero et al. 2021; García-Bernete et al. 2024a; Zhang et al. 2024)111https://gatos.myportfolio.com. The full GATOS sample is a nearly complete selection of AGN with luminosities at distances of Mpc from the 70 Month Swift-BAT All sky Hard X-ray Survey (Baumgartner et al. 2013). This sample is also largely unbiased to obscuration/absorption even up to column densities of , from which the six targets of GO program #1670 are further selected based on their outflow properties. A first analysis of the data set of this program, focussed on silicate features and water ice at small scales, is presented by García-Bernete et al. (2024a), followed by analyses focused on different aspects of one or more of the six targets (e.g., Davies et al. 2024; Esparza-Arredondo et al. 2024; García-Bernete et al. 2024c; Hermosa Muñoz et al. 2024; Zhang et al. 2024, and Delany et al. in prep., Haidar et al. in prep.).
@@
In this letter, we focus on the remaining three targets, which exhibit relatively weak PAH emission without significantly extended outflow regions and cover a considerable range of AGN strength. The goal of this work is to shed light on the specific mechanisms that could be responsible for the variation of PAH properties around different AGN. To this end, this letter combines the analysis of both observational measurements and theoretical models of not only PAH features but also other infrared emission lines, which is among the first attempts of such work in the era of JWST. Given the relatively weak PAH emission of the three targets studied here, after some testing work, we choose to proceed the analysis with the MRS spectra extracted with a series of apertures (see Figure 1, 500 pc in physical scale for each, well beyond the nuclear regions studied by García-Bernete et al. 2024c), rather than the spatially resolved MRS spectra. These apertures are simply selected to match the field of view of MRS IFU units.
@@ 3 PAH Characteristics of Central Apertures in Target Seyferts
Compared to the SF spaxels, the apertures of ESO137-G034 overall exhibit the much larger PAH 11.3 /7.7 band ratios and slightly larger PAH 6.2 /7.7 band ratios, while the aperture of MCG-05-23-016 exhibits the much smaller PAH 6.2 /7.7 band ratio and the moderate PAH 11.3 /7.7 band ratio. Moreover, the apertures of ESO137-G034 and MCG-05-23-016 on the PAH diagram are close to the loci of low metallicity star-forming systems (i.e., H II regions and BCDs) and low redshift quasars, respectively. We will specifically discuss the similarity and difference between these systems in Section 4 to shed more light on the physical mechanisms that could be responsible for their PAH characteristics. Meanwhile, the apertures of NGC 3081 exhibit the PAH band ratios of a wide distribution, overlapping with those of a large sample of Seyfert galaxies studied in the era of Spitzer. Overall, apertures of targets studied here have a wide distribution of PAH band ratios, covering those of most Seyferts, as well as those of the outflow regions and the innermost regions of the three targets studied in Paper I.
@@ 4 Emission Line Diagnostics of Central Apertures in Target Seyferts
We find that although the three apertures in ESO137-G034 exhibit the ionized emission line ratios consistent with the results obtained by AGN models, their highest ionization potential lines (e.g., [Ne V] with the ionization potential, i.e., IP, of 97.1 eV) are remarkably weak compared to other high ionization potential lines (e.g., [O IV] and [S IV] with the IP of 54.9 eV and 34.8 eV, respectively). For reference, the apertures in NGC 3081, which are also dominated by AGN excitation, exhibit overall strong emission of all the three high ionization potential lines. Specifically, AGN models calculated with incident radiation SEDs of lower Eddington ratios and smaller ionization parameters, i.e., as the indicator of ionization field intensity, exhibit lower [Ne V]/[Ne II] ratios (Pereira-Santaella et al. 2024). Accordingly, the lower [Ne V]/[Ne II] ratios reflect the weak accretion nature and moderately intense radiation field of ESO137-G034. Additionally, the radiation filed therein is hard given the large values of the [Ne III]/[Ne II] ratio, which is a canonical indicator of the radiation filed hardness (e.g., Thornley et al. 2000; Groves et al. 2006). Meanwhile, we find that the central region of MCG-05-23-016 is likely also affected by fast radiative shocks in addition to AGN excitation. However, the emission lines from the central aperture of MCG-05-23-016 do not show the significant broadening as expected for highly shocked regions. This suggests that the fast radiative shocks in MCG-05-23-016 are more likely playing a role in the radiative rather than kinetic manner as detailed later.
@@
Related to this result, we find that [Fe II]/Pf and [Fe II]/Hu ratios, as another empirical diagnostics of collisional shock heating, of the three targets exhibit the same trend as their ratios. Namely, the apertures of ESO137-G034 and MCG-05-23-016 exhibit the highest and lowest [Fe II]/Pf as well as [Fe II]/Hu ratios, respectively, while the apertures of NGC 3081 exhibits the values between those of ESO137-G034 and MCG-05-23-016 (see Figure A3). In fact, the intrinsic [Fe II] emission of MCG-05-23-016 is not weak, and its value is roughly the same as that of the innermost aperture of NGC 3081 and greater than those of other apertures of NGC 3081 (see Table A2). The aperture of MCG-05-23-016 exhibits the relatively weak evidence of collisional shock heating, but have the strongest [Ne II] emission and the highest Pf/ and Hu/ ratios among all the 13 apertures of the three targets (see Figure A2). Meanwhile, the three apertures of ESO137-G034 have the highest temperatures according to their ratios, consistent with its strongest signature of collisional shock heating. These results, combining the findings from Figure 4, indicate that the central region of MCG-05-23-016 is more likely dominated by photo-ionization from the (fast radiative) shock precursor in addition to the AGN, i.e., the radiative mode feedback, rather than the collisional shock excitation, i.e., the kinetic mode feedback, while the central region of ESO137-G034 is highly heated due to the collisional shock excitation (see further discussion in Section 5).
@@ 5 Discussion: Underlying Mechanisms Responsible for the PAH Characteristics
Accordingly, combining the PAH characteristics and other emission line diagnostics covered by JWST spectra, we are able to have a more comprehensive understanding of different feedback effects of AGN activity, which is pivotal for fully understanding the galaxy formation and evolution. Further study, similar to the analysis done here, of a larger sample and for different AGN systems (e.g., quasars, LINERs, and Seyferts) could allow us to use PAH characteristics as diagnostics of different evolutionary processes in galaxies, which has a promising application in the era of JWST.
@@ 6 Summary and Conclusions
Leveraging high quality JWST MIRI/MRS IFU observations, this letter showcases the distinct PAH characteristics in a sample of three Seyferts, based on the PAH measurements from a series of (i.e., 500 pc) apertures (Section 2). We find that positions occupied by apertures in the three targets on the PAH diagram are different from those occupied by regions in star-forming galaxies. Specifically, we find larger PAH 11.3 /7.7 ratios or smaller PAH 6.2 /7.7 ratios for apertures in the three targets, indicating overall more neutral or larger PAHs therein compared to regions in star-forming galaxies (Section 3). In addition to the PAH diagram, we also present and discuss other emission line diagnostics, revealing the existence of extra heating or potential fast radiative shocks in these targets (Section 4). Combining the PAH diagram and the emission line diagnostics, we further discuss the underlying mechanisms responsible for the PAH characteristics of the three targets (Section 5).